Message: A body having a mass of M starts from rest at the point x=0. It accelerates by -t^2+4t.
The maximum distance that is covered by the body in the +ve x direction before moving backward in the -ve direction is
1
Expert's answer
2020-07-09T10:08:37-0400

We have that


"a=-t^2+4t" .


So,


"v=\int{adt}=-\frac{t^3}{3}+4\frac{t^2}{2}"


and


"S=\int{vdt}=-\frac{t^4}{12}+2\frac{t^3}{3}"


"S=S_{max}" at "v=0"


"v=-\frac{t^3}{3}+4\frac{t^2}{2}=0\to t=6s" (if "t=0" then "v=0" )


"S=S_{max}=-\frac{t^4}{12}+2\frac{t^3}{3}=-\frac{6^4}{12}+2\frac{6^3}{3}=36m"